### v2

# MIRA: Medical Time Series Foundation Model for Real-World Health Data

Abstract
A unified foundation model for medical time series—pretrained on open access and ethics board-approved medical corpora—offers the potential to reduce annotation burdens, minimize model customization, and enable robust transfer across clinical institutions, modalities, and tasks, particularly in data-scarce or privacy-constrained environments. However, existing generalist time series foundation models struggle to handle medical time series data due to their inherent challenges, including irregular intervals, heterogeneous sampling rates, and frequent missing values. To address these challenges, we introduce MIRA, a unified foundation model specifically designed for medical time series forecasting. MIRA incorporates a Continuous-Time Rotary Positional Encoding that enables fine-grained modeling of variable time intervals, a frequency-specific mixture-of-experts layer that routes computation across latent frequency regimes to further promote temporal specialization, and a Continuous Dynamics Extrapolation Block based on Neural ODE that models the continuous trajectory of latent states, enabling accurate forecasting at arbitrary target timestamps. Pretrained on a large-scale and diverse medical corpus comprising over 454 billion time points collect from publicly available datasets, MIRA achieves reductions in forecasting errors by an average of 10% and 7% in out-of-distribution and in-distribution scenarios, respectively, when compared to other zero-shot and fine-tuned baselines. We also introduce a comprehensive benchmark spanning multiple downstream clinical tasks, establishing a foundation for future research in medical time series modeling.
footnotetext:  Work done during research internship at Microsoft Research.footnotetext:  Corresponding author: chanx@microsoft.com.
1 Introduction
Medical time series data, including signals such as electrocardiograms (ECG) [1], electroencephalograms (EEG) [2], vital signs and laboratory measurements [3], are key to understanding the dynamic physiological states of patients over time [4, 5]. Continuously modeling these patient data trajectories supports clinical forecasting tasks such as predicting organ failure or treatment response [1, 6], enabling earlier decisions and better patient outcomes [7, 8]. However, effectively utilizing medical time series data in practice remains challenging, as patient populations, disease profiles, and clinical protocols can vary widely across regions and institutions. Moreover, such data is often collected in an irregular and asynchronous manner [9, 10], further complicating efforts to build generalizable models. These challenges are compounded by regulatory constraints such as GDPR [11, 12], which inhibit data sharing and result in institutions training task-specific models from scratch on local data, incurring significant resource overhead. This highlights the need for a shareable medical foundation model trained on validated, open medical time series datasets, that is able to learn generalizable patterns from large and varied datasets, reducing the need for extensive data annotation and custom model building, while also helping to share knowledge effectively across different clinical institutions, data types, and medical tasks [13, 14].
Figure 1: Medical time series exhibit irregular intervals, heterogeneous sampling rates, and frequent missingness driven by clinical workflows.
However, creating effective medical foundation models for time series is challenging, because clinical data are inherently highly irregular and varied: signals like ECGs recorded at millisecond intervals [15] are found alongside laboratory tests conducted hours apart [16] (Figure 1). Differences in clinical workflows and equipment limits result in irregular sampling, missing values, and inconsistent temporal dependencies [17, 18]. Furthermore, physiological variables differ widely in their value and frequency ranges, requiring models that can handle these varied scales and types of data. Traditional methods using medical task-specific time-series models [19, 20, 21, 22, 23, 24], trained on isolated datasets, often do not generalize well to new situations and are difficult to scale [25, 26, 9]. Meanwhile recent general-purpose time series foundation models [27, 28, 29, 30, 31, 32] show promise, but they usually assume uniform time intervals. Similarly, emerging medical time series foundation models demonstrate cross-dataset generalization but only within narrowly defined domains—for example, across EEG datasets for sleep monitoring or across datasets for Alzheimer’s diagnosis, and cannot handle continuous-time forecasting or irregular sampling [33, 34, 35, 36]. These gaps—where current models struggle with the complex, irregular, and multi-frequency nature of medical data [27, 37]—motivates our research and raises a central question:
How can we design a scalable and generalizable foundation model that captures the irregularity and multi-frequency nature of medical time series, while enabling robust transfer across diverse medical tasks?
To address this central question, we introduce MIRA—a Medical Foundation Model for Irregularity Adaptation. MIRA is specifically designed to address the challenges of medical time series: First, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE), which extends the standard RoPE [38] to handle time values using learnable frequency adjustments, allowing MIRA to accurately model the varying frequency intervals of irregularly sampled data. Second, to further improve its ability to adapt to different time features in medical signals, MIRA uses a frequency-specific mixture-of-experts (MoE) layer that learns to route frequency-related information through specialized components. Finally, MIRA integrates a Continuous Dynamics Extrapolation Block, using Neural Ordinary Differential Equations (Neural ODEs) [39]. This feature allows the model to predict patient health paths at any future time point, moving beyond the limits of fixed observation grids. MIRA is pretrained on a large curated collection of medical time series, comprised of over 454 billion time points from publicly available datasets. These datasets cover readings from intensive care units (ICUs), operating rooms, pediatric critical care, and long-term sleep and mental health monitoring. All data undergo preprocessing that standardizes time alignment, normalizes sampling frequencies, and maintains clinical realism. MIRA outperforms other time series foundation models with a similar number of activated parameters across various real-world benchmarks, achieving reductions in forecasting errors by an average of 10% and 7% in out-of-distribution and in-distribution scenarios, respectively.
To summarize, this paper makes the following contributions: First, we propose MIRA, a new foundation model specially designed for medical time series forecasting. MIRA directly tackles the challenges of irregular sampling and varied temporal dynamics within a single, well-structured system. Second, we curate, preprocess and release an extensive and diverse corpus of medical time series, containing over 454 billion observations, to be used for model pre-training. This collection, built from multiple public datasets, is curated to cover diverse types of medical time series. We select data sources and control their proportions to reflect real-world variability, ensuring correct time alignment, consistent types, and high quality. Third, we establish a comprehensive benchmark suite that covers a wide range of clinical forecasting tasks. This suite allows for consistent evaluation of MIRA and future models, encouraging further research and development into robust and generalizable medical time series models.
2 Related Work
Medical Time Series Models. Machine learning has enabled significant progress in healthcare time series analysis, supporting a wide range of tasks, including dynamic forecasting [20, 21, 40, 41], survival analysis [42, 43], clustering and phenotyping [44, 45, 46], screening and monitoring [22, 47, 48], early diagnosis [23, 24, 49], pharmacology [50], treatment effect estimation [51, 52, 41], epidemiological and pandemic influenza surveillance [53], and hospital resource allocation [19, 54]. Despite this progress, most existing methods are narrowly designed for specific tasks or datasets, limiting their adaptability to diverse clinical scenarios [55]. Recently, foundation models for medical time series have emerged, showing promising cross-dataset generalization [33, 36, 34]. However, these models focus primarily on classification tasks and are not designed to handle continuous-time forecasting or irregular temporal patterns common in real-world clinical data.
Generalist Time Series Foundation Models. Recent foundation models have shown strong zero-shot forecasting without task-specific tuning [56, 57, 58, 59, 60]. Chronos [32] first adapted T5 [61] by discretizing time series intervals as text tokens. More recent models—Moirai [27], Sundial [31], and TimesFM [29]—improve support for variable dimensions and frequencies. Time-MoE [28] and Moirai-MoE [30] further improve specialization with sparse experts. However, these models largely require regularly sampled data, limiting their use on irregular clinical data. See Appendix A for more information.
Irregular Time Series Models. Modeling irregular time series has attracted increasing attention, particularly for applications requiring continuous-time reasoning. Early works such as Neural ODEs [39, 62] and State Space Models (SSMs)[63, 64] have demonstrated the ability to capture continuous dynamics by parameterizing the evolution of latent states over time. Another line of research focuses on adapting deep neural architectures, such as RNNs [65, 66, 67] and Transformers [18, 68, 69, 70], to irregular settings. While these models have shown success in tasks like interpolation and classification, their effectiveness for long-term forecasting remains underexplored.
3 Methodology
In this section, we discuss the architecture of MIRA. We first formalize the irregular medical time series forecasting task (Section 3.1). We then introduce the model architecture (Section 3.2). Finally, we present the pretraining corpus and training strategy (Section 3.3).
3.1 Irregular Medical Time Series Forecasting
Let each time series instance be represented as paired sequences of timestamps and observations: $\{(t_{i},x_{i})\}_{i=1}^{N}$ where $t_{i}\in\mathbb{R}^{+}$ are strictly increasing ($t_{1}<\cdots<t_{N}$) and $x_{i}\in\mathbb{R}\cup\{\text{NaN}\}$ contains real values or missing entries. Our model is designed to handle two prevalent irregular patterns: (1) Regular-grid missing values. Observations occur at equidistant timestamps $t_{i}=i\Delta t$ but with time-level missing values: some timestamps may have missing measurements. Formally, there exists a subset $\mathcal{M}\subseteq{1,\ldots,N}$ such that the corresponding observations $x_{i}$ are missing (i.e., $x_{i}=\text{NaN}$) for all $i\in\mathcal{M}$, while the timestamps ${t_{i}}$ remain fully observed. (2) Irregular sampling. Timestamps $t_{i}$ follow non-uniform intervals with $\Delta t_{i}=t_{i+1}-t_{i}$ varying for different $i$. All observations at existing timestamps are present ($x_{i}\in\mathbb{R}$), but the temporal spacing is irregular. Adhering to the channel-independent setting [59], we formulate the problem in a uni-variate time series manner [27]. Given history $\{(x_{i},t_{i})\}_{i=1}^{L}$, the forecasting task requires predicting future values at known target timestamps $\{t_{L+1},...,t_{L+H}\}$:
$\hat{\mathbf{x}}_{L+1:L+H}=f_{\theta}\big{(}\{x_{i},t_{i}\}_{i=1}^{L};\{t_{j}\}_{j=L+1}^{L+H}\big{)}\in\mathbb{R}^{H}$ (1)
where $\{t_{j}\}_{j=L+1}^{L+H}$ may follow either regularity pattern.
3.2 Model Overview
Figure 2: Architecture of MIRA. Takes irregular medical time series and timestamps as input, applying CT-RoPE for continuous temporal encoding. A Sparse Temporal Mixture-of-Experts layer routes tokens to specialized experts based on frequency. A Continuous Dynamics Extrapolation Block evolves latent states toward arbitrary target timestamps for flexible time-aware forecasting.
We present MIRA, a decoder-only architecture for universal irregular medical time series forecasting. As illustrated in Figure 2, MIRA consists of three key components: (1) a Continuous-Time Rotary Positional Encoding, (2) a Sparse Temporal Mixture-of-Experts Module, and (3) a Continuous Dynamics Extrapolation Block. Together, they enable scalable, frequency-adaptive modeling of non-uniform clinical sequences.
3.2.1 Continuous-Time Rotary Positional Encoding
Standard RoPE [38] assumes discrete and uniformly spaced token indices, which limits its applicability to clinical time series with real-valued and irregular timestamps. To address this, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE) that generalizes RoPE to operate directly on continuous-time inputs. Let $t\geq 0$ denote the timestamp of a token in an irregularly sampled sequence, and let $d$ be the model dimensionality, assumed to be even. Specifically, we discretize continuous timestamps $t\geq 0$ into rotation angles without assuming fixed intervals, enabling the model to handle irregular sampling (more detail can be found at Appendix B). The angular frequencies $\{\omega_{i}\}_{i=0}^{d/2-1}$ are defined as: $\omega_{i}=10000^{-2i/d}$. The resulting time-dependent rotation angle is given by:
$\theta_{i}(t)=\omega_{i}\cdot t$ (2)
Given an input embedding $\mathbf{x}\in\mathbb{R}^{d}$, we partition it into $d/2$ two-dimensional sub-vectors $(x_{2i},x_{2i+1})$, and apply a planar rotation using the computed angle $\theta_{i}(t)$:
$\begin{bmatrix}x^{\prime}_{2i}\\ x^{\prime}_{2i+1}\end{bmatrix}=\begin{bmatrix}\cos\theta_{i}(t)&-\sin\theta_{i}(t)\\ \sin\theta_{i}(t)&\cos\theta_{i}(t)\end{bmatrix}\begin{bmatrix}x_{2i}\\ x_{2i+1}\end{bmatrix}$ (3)
CT-RoPE with Attention. The design of the rotary position encoding naturally supports relative position modeling. Specifically, the inner product between rotated query and key vectors depends only on the difference between their timestamps. This can be seen by expanding the attention score:
$\langle q_{m},k_{n}\rangle=x_{m}^{\top}(W^{Q})^{\top}R_{\Theta,m}^{\top}R_{\Theta,n}W^{K}x_{n}$ (4)
Since both $R_{\Theta,m}$ and $R_{\Theta,n}$ are rotation matrices constructed from trigonometric functions of the timestamp-derived angles, their product is itself a rotation matrix dependent only on the time difference:
$R_{\Theta,m}^{\top}R_{\Theta,n}=R_{\Theta,n-m}$ (5)
This implies that the positional interaction between two tokens is solely a function of their relative timestamp offset $(n-m)$. In CT-RoPE, this property generalizes to real-valued timestamps, enabling the attention mechanism to capture continuous-time relational structure while retaining the efficiency and structure of standard dot-product attention. In line with Chowdhery et al. [71], we remove biases except in QKV projections to enhance extrapolation. More detail can be found in Appendix C.
3.2.2 Frequency-Specific Mixture-of-Experts Block
Medical time series often exhibit dynamics across multiple temporal frequencies, ranging from smooth long-term trends to rapid, short-term variations. To effectively capture such heterogeneity while maintaining computational efficiency, we adopt a sparse Mixture-of-Experts (MoE) architecture to replace the standard feedforward sub-layer in each Transformer block. In practice, each token is routed to a subset of $K$ experts selected from a shared pool of $N$ lightweight feedforward networks. These experts are parameterized independently and are intended to specialize in distinct temporal or semantic structures. Additionally, a shared expert is universally applied to all tokens, serving as a global residual pathway. Given the token representation $\mathbf{\bar{u}}^{l}_{t}\in\mathbb{R}^{D}$ at layer $l$, the output of the MoE block is calculated as:
$\operatorname{MoE}\left(\mathbf{\bar{u}}^{l}_{t}\right)$ $=g_{N+1,t}\cdot\mathrm{FFN}_{N+1}\left(\mathbf{\bar{u}}^{l}_{t}\right)+\sum_{i=1}^{N}g_{i,t}\cdot\mathrm{FFN}_{i}\left(\mathbf{\bar{u}}^{l}_{t}\right),$ (6)
where $\mathrm{FFN}_{i}(\cdot)$ denotes the $i$-th expert network, $\mathrm{FFN}_{N+1}(\cdot)$ the shared expert, and $g_{i,t}$, $g_{N+1,t}$ the corresponding routing weights. The non-shared expert weights $g_{i,t}$ are obtained via a softmax gating mechanism followed by top-$K$ selection:
$s_{i,t}$ $=\frac{\exp\left((\mathbf{W}_{i}^{l})^{\top}\mathbf{\bar{u}}^{l}_{t}\right)}{\sum_{j=1}^{N}\exp\left((\mathbf{W}_{j}^{l})^{\top}\mathbf{\bar{u}}^{l}_{t}\right)},\quad g_{i,t}$ =$=\begin{cases}s_{i,t},&\text{if }s_{i,t}\in\operatorname{TopK}(\{s_{j,t}\}_{j=1}^{N},K),\\ \\ 0,&\text{otherwise},\end{cases}$ (7)
where $\mathbf{W}_{i}^{l}\in\mathbb{R}^{D}$ are the trainable gating vectors. The shared expert weight $g_{N+1,t}$ is computed independently using a sigmoid gate:
$g_{N+1,t}=\sigma\left((\mathbf{W}_{N+1}^{l})^{\top}\mathbf{\bar{u}}^{l}_{t}\right),$ (8)
where $\sigma(\cdot)$ denotes the element-wise sigmoid function and $\mathbf{W}_{N+1}^{l}\in\mathbb{R}^{D}$ is the shared expert gating vector.
3.2.3 Continuous Dynamics Extrapolation Block
Auto-regressive transformer architectures, which generate predictions in a stepwise manner under causal masking, cannot incorporate the timestamp of the target token during inference, as it is not available until after generation. To address this limitation and enable extrapolation to arbitrary timestamps, we introduce a Neural ODE-based [39] extrapolation module that evolves the latent state from the current token’s timestamp to the target prediction token timestamp, allowing time-aware forecasting at unseen or irregular time points.
Given $h(t_{N})\in\mathbb{R}^{d}$ (the state at time $t_{N}$) and the next target timestamp $t_{N+1}$, the Neural ODE module extrapolates $h_{N}$ to $h(t_{N+1})$. We define the temporal evolution of the hidden state $h(s)$ over the interval $s\in[t_{N},t_{N+1}]$ as:
$\frac{dh(s)}{ds}=f(s-t_{N},h(s);\theta_{ODE}),\quad h(t_{N})=h_{N},\quad\text{for }s\in[t_{N},t_{N+1}]$ (9)
where $f:\mathbb{R}_{\geq 0}\times\mathbb{R}^{D_{model}}\to\mathbb{R}^{D_{model}}$ is the dynamics function, parameterized by a neural network (e.g., an MLP) with parameters $\theta_{ODE}$. $f$ takes the relative time $\Delta s=s-t_{N}$ and the current state $h(s)$ as input. The state at the target time $t_{N+1}$ is obtained by integrating the ODE dynamics:
[…]
This integral is computed numerically using an ODE solver (i.e. the Dormand-Prince (RK45) method). Let the result of this numerical integration be denoted as $h^{\prime}_{N+1}=h(t_{N+1})$. More detail is in Appendix D.
Implementation with Adaptive ODE Solvers We numerically approximate the solution to Equation 10 using adaptive step-size ODE solvers. Appropriate absolute and relative error tolerances (e.g., $10^{-6}$) are set to manage the trade-off between accuracy and computational cost.
3.3 Model Training
Medical Pretraining Dataset.
To support generalizable and clinically relevant time series modeling, we curate a large pretraining corpus spanning over 454 billion time points from various real-world healthcare settings. The dataset collection includes signals from ICUs a nd operating rooms, pediatric critical care, long-term sleep and mental health monitoring, and population-level epidemiological surveillance. All data are drawn from publicly available clinical datasets, including MIMIC-III [72], MIMIC-IV [73], PTB-XL [74], Sleep-EDF [75], and the WAVES Pediatric Waveform Database [76]. A full summary of included datasets and the pre-processing applied is provided in Appendix E.
Loss Function.
Training large-scale medical time series models requires balancing predictive accuracy, numerical stability, and sparse expert utilization. To ensure robustness against outliers and noisy measurements common in clinical data, we employ the Huber loss $\mathcal{L}_{\text{Huber}}$ over autoregressive multi-horizon predictions:
$\mathcal{L}_{\text{Huber}}(x,\hat{x})=\begin{cases}\frac{1}{2}(x-\hat{x})^{2},&\text{if }|x-\hat{x}|\leq\delta,\\ \delta\cdot(|x-\hat{x}|-\frac{1}{2}\delta),&\text{otherwise},\end{cases}$ (11)
where $\delta$ is a threshold controlling the transition between L2 and L1 loss regimes. To avoid expert collapse in the sparse MoE layer, we introduce a load balancing loss $\mathcal{L}_{\text{aux}}$ that promotes uniform usage of experts. Let $f_{i}$ be the fraction of tokens assigned to expert $i$, and $r_{i}$ the average routing probability:
$\mathcal{L}_{\text{aux}}=N\cdot\sum{i=1}^{N}f_{i}r_{i},$ (12)
where $f_{i}=\frac{1}{KT}\sum_{t=1}^{T}\mathbb{I}(\text{Expert}i\text{ selected at }t)$ and $r_{i}=\frac{1}{T}\sum{t=1}^{T}s_{i,t}$ with $s_{i,t}$ being the softmax routing score.
Model Configurations.
Table 1: A high-level summary of MIRA model configurations.
Layers Experts $K$ $d_{\mathrm{model}}$ $d_{\mathrm{ff}}$ $d_{\mathrm{expert}}$ Activated Params Total Params
$MIRA_{small}$ 8 8 2 288 1152 144 30 $\mathrm{M}$ 73 $\mathrm{M}$
$MIRA_{base}$ 12 8 2 384 1536 192 50 $\mathrm{M}$ 114 $\mathrm{M}$
$MIRA_{large}$ 12 8 2 768 3072 384 200 $\mathrm{M}$ 455 $\mathrm{M}$
We adopt a similar model configuration strategy following Time-MoE [28] by providing three model variants of increasing scale: $MIRA_{small}$, with approximately 73 million parameters; $MIRA_{base}$, with approximately 114 million parameters; $MIRA_{large}$, with approximately 455 million parameters;All models are trained on max to eight NVIDIA 80G H/A100 GPUs, using a micro batch size of 128 , and a maximum sequence length of 512. For model configurations and training details, refer to Appendix F.
4 Experiments
In this section, we present the empirical evaluation of MIRA. We begin by outlining the experimental setup (Section 4.1). We then evaluate MIRA on zero-shot forecasting benchmarks under both out-of-distribution (Section 4.2) and in-distribution (Section 4.3) settings. Additionally, we analyze the model’s scaling behavior across varying model and dataset sizes, and evaluate its robustness to different levels of data irregularity (Section 4). Finally, we conduct ablation studies to examine what contributes to MIRA’s performance (Section 4.5).
4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, pharmacological safety, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [77] and CinC 2019 [78]. . Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [79], Johns Hopkins COVID-19 Dataset [80] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, WHO fluent  [^2] https://www.who.int/tools/flunet, Heart Rate [81] and illness [82], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
Baselines. We compare MIRA against 13 state-of-the-art forecasting models, which we categorize into three groups: (i) zero-shot foundation models, including Time-MoE [28]  [^3] We excluded Time-MoE Ultra and Sundial from baselines as they are not open-sourced., Moirai [27], Moirai-MoE [30], Moment [83], TimeGPT [84], Timer [85], Lag-Llama [86], TimesFM [29], and Chronos [32], which require inputs to be interpolated for evaluation; and (ii) full-shot forecasting models, including ContiFormer [18], T-PatchGNN [87], Neural-CDE [88], and ODE-RNN [65], which are specialized for irregularly sampled time series and require task-specific training. (iii) Continue pre-trained zero-shot foundation models on medical corpora, including Time-MoE [28], Moirai [27] and Chronos [32] which performance best on zero-shot evaluation. Implementation details for all baselines are provided in Appendix H.
Evaluation Metrics. We measure the Root Mean Squared Error (RMSE) and the Mean Absolute Error (MAE) as the evaluation metrics. Detailed definitions are provided in Appendix I.
Table 2: Zero-shot forecasting performance on out-of-distribution datasets that are regularly sampled but contain missing values. Reported values are averaged across all prediction lengths. Lower RMSE and MAE indicate better predictions. Red: the best; Blue: the second best compared to zero-shot baselines; Underline: the best performance compared to full-shot baselines.
Models Zero-shot Ours Full-shot Time Series Models
MIRAsmall MIRAbase MIRAlarge Contiformer T-PatchGNN ODE-RNN Neural-CDE TimesFM
Metrics RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Cinc 2012 ($10^{1}$) $5.963$ $5.891$ $5.764$ $5.732$ $5.312$ $5.120$ $5.987$ $5.985$ $6.247$ $6.246$ $6.997$ $6.995$ $7.498$ $7.497$
Cinc 2019 ($10^{1}$) $7.456$ $6.597$ $7.347$ $7.311$ $7.212$ $7.246$ $6.989$ $7.072$ $7.608$ $7.541$ $10.293$ $9.7724$ $9.591$ $8.322$
[…]
MIT-BIH $0.293$ $0.198$ $0.163$ $0.111$ $0.122$ $0.091$ $0.453$ $0.354$ $0.705$ $0.627$ $0.882$ $0.623$ $0.242$ $0.196$ $0.335$ $0.141$
CDC-IHA ($10^{1}$) $5.976$ $4.684$ $5.729$ $4.502$ $5.534$ $4.401$ $5.510$ $4.133$ $9.522$ $7.974$ $10.068$ $9.052$ $7.892$ $6.766$ $15.633$ $4.408$
JH COVID-19 ($10^{2}$) $0.407$ $0.355$ $0.504$ $0.349$ $0.478$ $0.336$ $0.323$ $0.297$ $0.350$ $0.291$ $0.424$ $0.331$ $0.545$ $0.503$ $2.329$ $0.322$
ILI $1.294$ $1.077$ $1.218$ $1.024$ $1.114$ $0.926$ $0.391$ $0.224$ $0.195$ $0.143$ $0.410$ $0.264$ $0.423$ $0.314$ $2.034$ $1.333$
[…]
Zero-shot Foundation Models Pre-trained on General-domain Corpora
Baseline Lag-Llama TimeGPT Timer Momentsmall Momentbase Momentlarge Moirai-MoEsmall Moirai-MoEbase
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) $1.764$ $1.488$ $2.258$ $1.915$ $1.901$ $1.704$ $2.966$ $1.852$ $2.939$ $1.835$ $2.917$ $1.735$ $1.982$ $1.600$ $2.144$ $1.652$
MIT-BIH $0.217$ $0.169$ $0.231$ $0.185$ $0.255$ $0.213$ $0.417$ $0.229$ $0.416$ $0.228$ $0.413$ $0.224$ $0.208$ $0.167$ $0.208$ $0.167$
CDC-IHA ($10^{1}$) $6.531$ $4.846$ $6.654$ $4.860$ $6.424$ $4.857$ $17.803$ $5.307$ $17.631$ $5.288$ $17.689$ $5.260$ $7.099$ $5.405$ $7.639$ $5.862$
JH COVID-19 ($10^{2}$) $3.596$ $1.432$ $1.879$ $1.580$ $2.647$ $2.328$ $3.077$ $0.553$ $3.097$ $0.554$ $3.064$ $0.549$ $1.452$ $0.725$ $19.391$ $3.190$
ILI $1.780$ $1.366$ $2.011$ $1.077$ $1.882$ $1.492$ $1.570$ $1.056$ $1.566$ $1.054$ $1.566$ $1.052$ $2.001$ $1.678$ $1.983$ $1.664$
[…]
Zero-shot Foundation Models Pre-trained on General-domain Corpora
Baseline Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) $2.359$ $1.965$ $2.156$ $1.767$ $2.098$ $1.644$ $0.850$ $0.650$ $0.833$ $0.639$ $1.357$ $0.587$ $1.189$ $0.489$ $1.218$ $0.506$
MIT-BIH $0.343$ $0.249$ $0.421$ $0.302$ $0.593$ $0.149$ $0.171$ $0.135$ $0.172$ $0.135$ $0.353$ $0.147$ $0.361$ $0.149$ $0.350$ $0.147$
CDC-IHA ($10^{1}$) $6.835$ $5.271$ $7.328$ $5.526$ $6.788$ $5.302$ $6.311$ $4.748$ $6.312$ $4.715$ $15.502$ $4.421$ $15.825$ $4.438$ $15.986$ $4.517$
JH COVID-19 ($10^{2}$) $1.917$ $0.695$ $0.991$ $0.474$ $0.614$ $0.402$ $0.596$ $0.402$ $0.512$ $0.371$ $4.826$ $1.031$ $3.835$ $0.551$ $3.478$ $0.521$
ILI $1.995$ $1.671$ $1.871$ $1.561$ $1.808$ $1.499$ $1.288$ $0.951$ $1.366$ $1.015$ $2.054$ $1.400$ $1.940$ $1.308$ $1.870$ $1.252$
[…]
Zero-shot Foundation Models Continue Pre-trained on Medical Corpora
Baseline Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) $2.047$ $1.685$ $1.907$ $1.536$ $1.601$ $1.263$ $0.648$ $0.524$ $0.603$ $0.488$ $1.049$ $0.555$ $0.965$ $0.488$ $0.902$ $0.458$
MIT-BIH $0.239$ $0.204$ $0.274$ $0.190$ $0.219$ $0.176$ $0.132$ $0.103$ $0.131$ $0.101$ $0.311$ $0.137$ $0.320$ $0.139$ $0.306$ $0.127$
CDC-IHA ($10^{1}$) $6.690$ $5.310$ $6.934$ $5.376$ $6.696$ $5.114$ $6.327$ $4.698$ $6.299$ $4.666$ $14.502$ $4.321$ $14.825$ $4.338$ $14.986$ $4.417$
JH COVID-19 ($10^{2}$) $0.579$ $0.448$ $0.619$ $0.478$ $0.812$ $0.337$ $0.509$ $0.353$ $0.517$ $0.362$ $4.225$ $0.947$ $3.535$ $0.510$ $3.328$ $0.469$
ILI $1.528$ $1.289$ $1.435$ $1.191$ $1.501$ $1.229$ $1.188$ $0.915$ $1.201$ $0.927$ $1.705$ $1.127$ $1.639$ $1.081$ $1.547$ $1.028$
[…]
4.2 Performance on Out-of-distribution Forecasting
Objective. We evaluated seven unseen benchmarks excluded from pre-training corpora. For comparison, we fine-tuned the full-shot forecasting models on the training split of each benchmark, while the zero-shot foundation models were evaluated directly without any task-specific training or fine-tuning. To validate the effectiveness of our model architecture, we additionally pre-trained existing foundation models on the same medical corpus. Notably, CINC 2012 and CINC 2019 were excluded from this evaluation for foundation models, as applying regular-grid interpolation at the finest resolution results in over 98% of time steps being interpolated, leading to poor performance for all baselines.
Results. Detailed results of out-of-distribution performance are reported in Table 2. First, MIRA consistently achieves state-of-the-art performance, outperforming all general-domain foundation models and specialized time series baselines. Specifically, MIRAlarge achieves the best results on all datasets, improving RMSE by over 10% on average compared to the strongest baselines, confirming the advantage of scaling and medical pretraining. This advantage is particularly pronounced on clinical benchmarks such as MIT-BIH and CDC-IHA, where MIRAlarge achieves both the lowest RMSE and MAE. Second, domain-specific pretraining proves essential. All model variants—continue pretrained on medical corpora—consistently outperform models trained on general time series data. This demonstrates the benefit of leveraging medical-specific temporal structures and distributions during pretraining. Interestingly, even smaller MIRA models surpass larger general-domain models, suggesting that data relevance is more critical than model size alone. Third, MIRA achieves performance close to or even exceeding fine-tuned full-shot models in several cases. For instance, $MIRA_{large}$ outperform than full-shot baseline on Cinc 2012 and slight lower in JH Covid-19.
4.3 Performance on In-distribution Forecasting
Table 3: Zero-shot forecasting performance on out-of-distribution datasets that are regularly sampled but contain missing values. Reported values are averaged across all prediction lengths. Lower RMSE and MAE indicate better predictions. Red: the best; Blue: the second best.
Models Zero-shot Ours Zero-shot Baseline
MIRAsmall MIRAbase MIRAlarge Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
RMSE SleepEDF ($10^{2}$) $0.215$ $0.195$ $0.189$ $0.301$ $0.668$ $0.304$ $0.228$ $0.244$ $0.411$ $0.414$ $0.413$
PTB-XL $0.147$ $0.127$ $0.121$ $0.177$ $0.270$ $0.416$ $0.110$ $0.109$ $0.228$ $0.234$ $0.229$
MIMIC-III $0.126$ $0.107$ $0.102$ $0.163$ $0.256$ $0.172$ $0.105$ $0.103$ $0.153$ $0.154$ $0.151$
MIMIC-IV $0.111$ $0.091$ $0.081$ $0.259$ $0.300$ $0.319$ $0.084$ $0.082$ $0.309$ $0.317$ $0.319$
WAVES $0.154$ $0.136$ $0.129$ $0.177$ $0.190$ $0.169$ $0.148$ $0.141$ $0.184$ $0.183$ $0.182$
[…]
MAE SleepEDF ($10^{2}$) $0.180$ $0.162$ $0.156$ $0.264$ $0.227$ $0.323$ $0.191$ $0.203$ $0.192$ $0.193$ $0.193$
PTB-XL $0.110$ $0.095$ $0.091$ $0.125$ $0.098$ $0.099$ $0.063$ $0.066$ $0.100$ $0.104$ $0.103$
MIMIC-III $0.106$ $0.089$ $0.084$ $0.141$ $0.164$ $0.138$ $0.081$ $0.078$ $0.079$ $0.080$ $0.080$
MIMIC-IV $0.094$ $0.069$ $0.061$ $0.223$ $0.291$ $0.143$ $0.064$ $0.062$ $0.134$ $0.142$ $0.143$
WAVES $0.129$ $0.112$ $0.106$ $0.157$ $0.181$ $0.155$ $0.124$ $0.116$ $0.119$ $0.118$ $0.117$
[…]
Objective. We evaluated in-distribution performance by holding out a portion of the pre-training datasets as test sets, ensuring no data leakage. All models were tested in zero-shot settings.
Results. As shown in Table 3, MIRA consistently achieves highly competitive zero-shot performance across all five pre-training datasets. Compared to baselines such as Moirai, Time-MoE, and Chronos, MIRA demonstrates lower RMSE and MAE on most datasets, particularly excelling on PTB-XL, MIMIC-III, and MIMIC-IV. This advantage holds across all model scales, indicating stable scalability and generalization. In contrast, baselines show larger variance or degradation on challenging datasets with missing values. These results validate the robustness of MIRA’s architecture in handling imperfect medical time series data without requiring task-specific adaptation.
4.4 Model Scaling and Data Behavior.
Table 4: Zero-shot forecasting performance on different missing rates. Reported values are averaged across all prediction lengths. Lower RMSE and MAE indicate better predictions. Red: the best; Blue: the second best.
Missing Rate 10%+ 20%+ 30%+ 40%+ 50%+ 60%+ 70%+ 80%+ 90%+ $1^{\text{st}}$ Count
Metrics ($10^{2}$) RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Ours MIRAsmall $3.15$ $2.82$ $3.28$ $2.95$ $3.41$ $3.08$ $3.54$ $3.20$ $3.67$ $3.34$ $3.80$ $3.47$ $3.93$ $3.61$ $4.07$ $3.75$ $4.22$ $3.90$ 00 00
MIRAbase $2.98$ $2.68$ $3.11$ $2.81$ $3.25$ $2.94$ $3.38$ $3.07$ $3.51$ $3.20$ $3.64$ $3.34$ $3.78$ $3.48$ $3.93$ $3.63$ $4.10$ $3.79$ 00 $1$
MIRAlarge $2.85$ $2.56$ $2.95$ $2.65$ $3.08$ $2.78$ $3.21$ $2.90$ $3.35$ $3.04$ $3.49$ $3.18$ $3.63$ $3.31$ $3.99$ $3.86$ $3.97$ $3.64$ $8$ $8$
[…]
Time-MoElarge $3.20$ $2.91$ $3.33$ $3.04$ $3.47$ $3.17$ $3.60$ $3.29$ $3.73$ $3.41$ $3.86$ $3.54$ $3.99$ $3.67$ $4.13$ $3.81$ $4.28$ $3.96$ 00 00
Moirailarge $4.80$ $4.18$ $5.39$ $4.64$ $5.68$ $4.55$ $6.02$ $4.83$ $6.09$ $5.06$ $6.18$ $5.02$ $6.31$ $5.09$ $6.51$ $5.24$ $6.71$ $5.36$ 00 00
Moiraibase $4.84$ $4.15$ $4.93$ $4.10$ $5.20$ $4.23$ $5.64$ $4.53$ $5.78$ $4.61$ $5.80$ $4.60$ $5.95$ $4.97$ $5.95$ $4.73$ $6.06$ $4.87$ 00 00
Moiraismall $4.91$ $4.22$ $5.04$ $4.20$ $6.12$ $4.93$ $6.23$ $4.98$ $6.44$ $5.49$ $6.47$ $5.18$ $6.59$ $5.60$ $6.66$ $5.20$ $6.86$ $5.71$ 00 00
Chronoslarge $4.71$ $4.10$ $4.85$ $4.22$ $5.02$ $4.31$ $5.31$ $4.52$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 00 00
Chronosbase $5.02$ $4.40$ $5.21$ $4.51$ $5.43$ $4.68$ $5.62$ $4.81$ $5.85$ $4.93$ $6.02$ $5.12$ $6.21$ $5.31$ $6.40$ $5.49$ $6.67$ $5.62$ 00 00
Chronossmall $5.32$ $4.71$ $5.49$ $4.82$ $5.68$ $4.93$ $5.81$ $5.02$ $6.01$ $5.20$ $6.21$ $5.32$ $6.40$ $5.49$ $6.58$ $5.61$ $6.83$ $5.72$ 00 00
[…]
Time-MoEbase $3.05$ $2.78$ $3.18$ $2.91$ $3.32$ $3.04$ $3.45$ $3.17$ $3.58$ $3.30$ $3.71$ $3.43$ $3.85$ $3.57$ $3.99$ $3.71$ $4.15$ $3.86$ 00 00
Moirailarge $4.41$ $3.76$ $4.92$ $4.20$ $5.35$ $4.41$ $5.69$ $4.76$ $5.93$ $4.80$ $6.09$ $4.92$ $6.23$ $4.98$ $6.48$ $5.29$ $6.62$ $5.30$ 00 00
Moiraibase $4.55$ $3.92$ $4.78$ $4.07$ $5.18$ $4.26$ $5.49$ $4.46$ $5.78$ $4.64$ $5.70$ $4.51$ $5.87$ $4.97$ $5.92$ $4.75$ $5.96$ $4.71$ 00 00
Moiraismall $5.02$ $4.21$ $5.36$ $4.40$ $5.55$ $4.98$ $5.97$ $4.92$ $6.25$ $5.34$ $6.43$ $5.11$ $6.44$ $5.35$ $6.57$ $5.41$ $6.58$ $5.30$ 00 00
Chronoslarge $4.30$ $3.89$ $4.58$ $4.02$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.61$ $4.73$ $5.83$ $4.91$ $6.02$ $5.10$ $6.31$ $5.32$ 00 00
Chronosbase $4.61$ $4.10$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 00 00
Chronossmall $4.83$ $4.31$ $5.02$ $4.43$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ $6.71$ $5.61$ 00 00
Objective. We investigate the robustness of the MIRA by comparing performance with baselines under varying missing datarates on the WHO FluNet dataset which is not used in pre-training.
Data Behavior Result. Table 4 summarizes the zero-shot forecasting performance under varying missing rates, ranging from 10% to 90%. MIRA consistently outperforms all baselines across all missing rates, with $MIRA_{large}$ achieving the lowest RMSE and MAE in nearly every setting. Notably, $MIRA_{large}$ maintains strong performance even as the missing rate increases, showing minimal performance degradation compared to other models. This demonstrates its robustness to severe information loss. While $Time-MoE_{large}$ performs competitively at lower missing rates, its performance drops more quickly as missingness increases.
Pretraining Scalability. Beyond downstream performance, we delve into the utilization of model capacity. We plot training curves in Figure 3. Compared with small size, the large version leads to totally 11% reduction in the converged training objective, showing significant performance promotion on in-distribution medical time series forecasting.
4.5 Ablation Analysis
Figure 3: (Left) Training curves on Medical corpora of different MIRA model sizes. (Right) Gating scores for experts across different layers in the there different frequence datasets.
Objective. We further perform an ablation study by removing key components of MIRA, to quantify their contribution, and the impact of the number of experts on performance.
Component Ablation Results. As shown in the left-hand side of Table 5, removing CT-RoPE (w/o CT-RoPE) leads to a performance drop, confirming the importance of our continuous-time positional encoding. Similarly, eliminating the Mixture-of-Experts block (w/o MoE Block) slightly degrades performance, showing the value of expert specialization. The largest degradation is observed when disabling the CT-Extrapolation Block, with RMSE and MAE increasing by 5 and 8%, respectively, highlighting its role in improving extrapolation. These results showcase the complementary benefits of all three components in achieving robust performance.
Table 5: Ablation studies. (Left) Evaluated with different model components. (Right) Analysis performance and inference speed across different topk setups. Lower values indicate better performance.
Component RMSE MAE
$MIRA_{base}$ 0.154 0.118
      w/o CT-RoPE $0.158$ $0.125$
      w/o MoE Block $0.157$ $0.122$
      w/o CT-Extrapolation Block $0.162$ $0.128$
Topk Setup RMSE MAE Speed (s/iter)
w/ {Top1} $0.160$ $0.121$ $0.097$
w/ {Top2} $0.154$ $0.118$ $0.101$
w/ {Top4} $0.154$ $0.117$ $0.112$
w/ {Top6} $0.156$ $0.120$ $0.124$
w/ {Top8} $0.159$ $0.122$ $0.127$
Expert Activation Analysis. We further examine the trade-off between predictive performance and inference efficiency by varying the number of activated experts ($top_{k}$), as shown in the right-hand side of Table 5. Activating two experts ($top_{2}$) achieves the best balance, reaching the lowest RMSE of 0.154 with an inference speed of 0.101 s/iter. Increasing to four experts provides no meaningful improvement in performance but increases inference time to 0.112 s/iter. Further increasing to six or eight experts leads to slower inference with diminishing returns. Conversely, using a single expert ($top_{1}$) sacrifices accuracy (0.160) for only a marginal speed gain. These findings indicate that using two experts offers an optimal trade-off between efficiency and accuracy for scalable deployment.
Activation Visualization.
We further visualize expert activation patterns on three datasets with distinct temporal resolutions: MIT-BIH (high-frequency, Hz-level), CDC-IHA (weekly), and COVID-19 (daily). As shown in Figure 3, low-frequency datasets such as Covid19 dataset tend to activate a different set of experts compared to the high-frequency MIT-BIH dataset.
5 Discussion & Conclusion
We introduce MIRA, a foundation model designed for medical time series forecasting under irregular conditions. By integrating CT-RoPE, a Time-Specialized MOE module, and Continuous Dynamics Extrapolation, MIRA demonstrates strong generalization capabilities across diverse medical datasets. This work highlights the potential of scalable and temporally adaptive solutions to real-world medical challenges.
Limitation. This work is based on publicly available, de-identified medical datasets. While these resources offer broad coverage and ensure reproducibility, they may not fully reflect the complexities of real-world clinical deployment. In addition, although these datasets are anonymized, residual privacy risks may still exist; however, addressing such risks lies beyond the scope of this work.
Appendix A Further Related Work
In this section, we delve deeper into the related work on (medical) time series foundation models. Current research efforts in universal forecasting with time series foundation models can be broadly classified into four categories, as summarized in Table 6:
(i) Encoder-only models. Moirai is trained on the LOTSA dataset comprising 27B time points, with model sizes up to 311M parameters [27]. MOMENT, based on the T5 architecture, is pretrained on the Time-series Pile dataset containing approximately 1B time points, reaching up to 385M parameters [83].
(ii) Encoder-decoder models, exemplified by Chronos [32], which tokenizes time series data via scaling and quantization, training on both public and synthetic datasets,offers pre-trained models at multiple scales, with the largest containing up to 710M parameters.
(iii) Decoder-only models. TimesFM is trained on a corpus of 100B time points, with model sizes up to 500M parameters [29]. Lag-Llama focuses on univariate probabilistic forecasting, utilizing a decoder-only Transformer architecture with up to 200M parameters [86]. Timer is a generative pre-trained Transformer model designed for large-scale time series modeling, with a base version containing 84M parameters and pre-trained on 260B time points [85].
(iv) Mixture-of-Experts architectures. Recent models adopt sparse Mixture-of-Experts (MoE) architectures to enhance scalability and efficiency. Time-MoE [28] scales to 2.4B parameters with only a few experts activated per input, while Moirai-MoE [30] achieves token-level specialization without frequency heuristics, improving adaptability and inference cost.
Table 6: Comparison between time series models.
Method MIRA Sundial Time-MoE Moirai-MoE Moirai TimesFM Moment Chronos Timer Lag-Llama TimeGPT
(Ours) (2025) (2024) (2024a) (2024) (2024) (2024) (2024) (2024d) (2023) (2023)
Architecture Decoder Decoder Decoder Decoder Encoder Decoder Encoder EncDec Decoder Decoder EncDec
(Max) Model Size 455M  [^4] M 2.4B 1.1B 311M 500M 385M 710M 67M 200M Unknown
Input Token Point Patch Point Patch Patch Patch Patch Point Patch Point Patch
Dataset Scale 454B 1TB 309B 27B 27B/231B* 100B 1.13B 84B 29B 0.36B 100B
Max Length 512 2880 4096 5000 5000 512 512 512 1440 1024 Unknown
FFN Sparse Dense Sparse Sparse Dense Dense Dense Dense Dense Dense Dense
* Depend on the way of calculation according to the original paper.
Appendix B Time Quantization Procedure
To ensure numerical stability and strictly increasing temporal inputs, we adopt a robust quantization strategy for irregular timestamps. The goal is to discretize time points with a suitable resolution while guaranteeing uniqueness. This is particularly important when encoding timestamps for attention-based models, where duplicate or non-monotonic inputs can lead to numerical or semantic inconsistencies.
The procedure begins with a predefined initial resolution (e.g., 1.0) and iteratively reduces it by a shrink factor (e.g., 10) until all quantized time points become unique. In each iteration, the algorithm rounds the original timestamps to the current resolution and checks for duplicates. If uniqueness is achieved, the quantized sequence is returned.
If, after a fixed number of iterations, uniqueness cannot be guaranteed, a fallback mechanism is triggered. In this case, small additive jitter is applied to break ties while preserving temporal ordering. The jitter is dynamically determined based on the magnitude of the input values, ensuring that the perturbation is both numerically safe and practically negligible.
Finally, the adjusted timestamps are cast to 32-bit floating-point precision and verified again for uniqueness. If rare collisions still exist due to floating-point rounding, a second corrective jitter step is applied.
This method ensures that all quantized timestamps are:
  • Strictly increasing, preserving temporal semantics;
  • Resolution-adaptive, automatically selecting the finest possible scale without manual tuning;
  • Numerically stable, avoiding floating-point collisions and degenerate attention behavior.
Appendix C Mathematical Analysis of CT-RoPE
C.1 Theoretical Properties
Linear Angle Scaling
For any timestamp $t\geq 0$, the rotation angles grow linearly with time:
$\theta_{i}(t)=\omega_{i}\cdot t$ (13)
where $\omega_{i}=10000^{-2i/d}$ are fixed angular frequencies.
Proof. Direct from Equation (1) in the main text. The linear formulation preserves temporal causality while maintaining bounded rotation magnitudes through exponentially decaying frequencies:
$\omega_{i}=\exp\left(-\frac{2i}{d}\ln 10000\right)$ (14)
This ensures higher dimensions receive progressively smaller rotations.
Relative Position Encoding
The inner product between rotated vectors depends only on the temporal difference:
$\langle R_{\Theta,t_{i}}q,R_{\Theta,t_{j}}k\rangle=\langle q,R_{\Theta,|t_{i}-t_{j}|}k\rangle$ (15)
Proof. Using rotation matrix properties:
$R_{\Theta,t_{i}}^{\top}R_{\Theta,t_{j}}$ $=\prod_{k=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{k}(t_{j}-t_{i}))&\sin(\omega_{k}(t_{j}-t_{i}))\\ -\sin(\omega_{k}(t_{j}-t_{i}))&\cos(\omega_{k}(t_{j}-t_{i}))\end{bmatrix}$ (16)
$=R_{\Theta,t_{j}-t_{i}}=R_{\Theta,|t_{i}-t_{j}|}$ (17)
The absolute difference emerges from $\cos(-\theta)=\cos(\theta)$ and $\sin(-\theta)=-\sin(\theta)$.
C.2 Rotation Matrix Construction
For a $d$-dimensional model (even dimension), define block-diagonal rotation matrix:
$R_{\Theta,t}=\bigoplus_{i=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{i}t)&-\sin(\omega_{i}t)\\ \sin(\omega_{i}t)&\cos(\omega_{i}t)\end{bmatrix}$ (18)
with key properties:
  • Norm Preservation: $\|R_{\Theta,t}x\|=\|x\|,\ \forall x\in\mathbb{R}^{d}$
  • Temporal Ordering: $t_{1}<t_{2}\Rightarrow\theta_{i}(t_{1})<\theta_{i}(t_{2})$
  • Differentiable: $\frac{\partial R_{\Theta,t}}{\partial t}$ exists $\forall t>0$
C.3 Attention Mechanism Extension
The scaled dot-product attention becomes:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{(R_{\Theta,t_{Q}}Q)(R_{\Theta,t_{K}}K)^{\top}}{\sqrt{d}}\right)V$ (19)
where $t_{Q}$ and $t_{K}$ are query/key timestamps respectively. This maintains:
  • Temporal Locality: $\lim_{\Delta t\to 0}\|R_{\Theta,t+\Delta t}-R_{\Theta,t}\|=\mathcal{O}(\Delta t)$
  • Causality: For $t_{Q}>t_{K}$, relative rotations prevent information leakage
  • Efficiency: Requires only $\mathcal{O}(d)$ extra computation vs standard attention
Appendix D Neural ODE
Gradient Computation via Adjoint Sensitivity
For end-to-end training, gradients of a loss function $\mathcal{L}$ (computed based on the final prediction) with respect to the ODE parameters $\theta_{ODE}$ and the initial state $h_{N}$ are efficiently computed using the adjoint sensitivity method [39]. This involves solving a backward-in-time ODE for the adjoint state $a(s)=\partial\mathcal{L}/\partial h(s)$:
$\frac{da(s)}{ds}=-a(s)^{\top}\frac{\partial f_{ODE}}{\partial h}(s-t_{N},h(s);\theta_{ODE})$ (20)
with the terminal condition $a(t_{N+1})=\partial\mathcal{L}/\partial h^{\prime}_{N+1}$. The required gradients are then computed via integrals involving the adjoint state $a(s)$.
Stability Considerations
To promote numerical stability and ensure the dynamics function $f_{ODE}$ remains well-behaved (e.g., satisfies Lipschitz conditions), techniques like spectral normalization can be applied to the weights $\theta_{ODE}$ of the neural network implementing $f_{ODE}$. This helps control the Lipschitz constant $L$ of $f_{ODE}$ with respect to $h$, contributing to bounded error propagation during integration as characterized by bounds of the form:
$\|h_{ODE}(t_{N+1})-h_{\text{true}}(t_{N+1})\|\leq\mathcal{O}\left(\frac{M}{L}(e^{L(t_{N+1}-t_{N})}-1)\right)$ (21)
where $M$ relates to the local truncation error of the numerical solver.
Algorithm 1 Neural ODE State Transition
0:  Hidden state $\mathbf{h}_{N}$, timestamps $t_{N}$, $t_{N+1}$
0:  Evolved state $\mathbf{\tilde{h}}(t_{N+1})$
1:  Configure ODE solver tolerances: $\texttt{rtol}\leftarrow 10^{-6}$, $\texttt{atol}\leftarrow 10^{-6}$
2:  Solve ODE: $\mathbf{\tilde{h}}(t_{N+1})\leftarrow\texttt{odeint}(f_{\text{ODE}},\mathbf{h}_{N},[t_{N},t_{N+1}])$
3:  Return final state: $\mathbf{\tilde{h}}(t_{N+1})[-1]$
D.1 Theoretical Guarantees
Our design satisfies the Picard-Lindelöf theorem conditions:
Theorem 1.
Let $f_{\text{ODE}}$ be spectrally normalized with maximum singular value $\sigma_{\max}$. Then for prediction horizon $\Delta t=t_{N+1}-t_{N}$, the state evolution admits a unique solution satisfying:
$\|\mathbf{\tilde{h}}(t_{N+1})-\mathbf{\tilde{h}}_{N}\|\leq\sigma_{\max}\Delta te^{\sigma_{\max}\Delta t}$ (22)
This bounded evolution property ensures numerical stability during long-horizon predictions.
Appendix E Pretraining Datasets
Table 7: The pre-training dataset of MIRA, which encompasses various sources.
Source WAVES MIMIC-III MIMIC-IV Sleep-EDF PTB-XL Total
(2023) (2016) (2020) (2000) (2020)
# Pts. 4.8B 400B 48B 0.14B 1.3B 454B
% 1.06% 88.06% 10.57% 0.03% 0.29% 100.0%
Our pretraining corpus consists of five publicly available medical time series datasets, selected for their diversity in sampling frequency, modality coverage, and clinical relevance:
MIMIC-III [72] is a widely used publicly available database containing de-identified health records of over 40,000 patients admitted to intensive care units (ICUs) at the Beth Israel Deaconess Medical Center between 2001 and 2012. It includes multivariate time series of vital signs, laboratory test results, medication administrations, and clinical notes. The data exhibit irregular sampling patterns and substantial missingness, making them suitable for evaluating temporal robustness.
MIMIC-IV [73] is the successor to MIMIC-III, covering ICU and emergency department admissions from 2008 to 2019. It offers expanded variable coverage, improved data standardization, and higher temporal resolution. Time series are recorded with greater granularity across a broader range of hospital units, including both adult and pediatric care. Like MIMIC-III, it reflects realistic clinical irregularity but with enhanced data fidelity and scale.
PTB-XL [74] is a 12-lead ECG dataset sampled at 100 Hz, containing over 20,000 clinical recordings annotated with diagnostic statements. It enables robust modeling of waveform-based cardiovascular signals.
Sleep-EDF [75] contains overnight sleep EEG and respiration recordings annotated with sleep stages, collected under controlled conditions. The data are sampled at 100 Hz and exhibit moderate regularity with biologically driven transitions.
WAVES [76] is a pediatric waveform dataset that combines high-frequency physiological signals (ECG, PPG, respiration) from intensive care and operating room settings. It provides long continuous sequences with natural variability and noise.
Appendix F Model Configurations and Training Detail
Informed by recent scaling studies [28], we design MIRA in three model scales: $MIRA_{small}$ (73M parameters), $MIRA_{base}$ (114M), and $MIRA_{large}$ (455M), enabling flexible deployment across different hardware constraints. All models are trained on up to eight NVIDIA 80GB A100 GPUs with a micro-batch size of 128 and a maximum sequence length of 512. We pre-trained one epoch with each training step processes approximately 65,000 time points. We consider forecast horizons of ${24,32,48,64}$ for short-term and long-term evaluation. Following standard practice, we apply an auxiliary load balancing loss with weight $\alpha=0.02$ to encourage expert utilization.
Appendix G Downstream Tasks and Datasets
Table 8: Characteristics of the WHO FluNet Dataset.
Dataset Time Points Missing Rate Frequency Irregular Sampling Signal Type
WHO FluNet 5.27M 61.52% Weekly Case Counts
CINC 2012 5.28M 0% Variable-specific ECG, Blood Pressure (ABP/NIBP), Respiration Rate, SpO2, Laboratory Measurements, etc.
MIT-BIH 31.1M 0% 360 Hz ECG
CinC 2019 39.5M 68.28% Hourly Heart Rate, Respiration Rate, Temperature, Blood Pressure
Johns Hopkins COVID-19 Dataset 3.81M 0% Daily Case Counts, Death Counts
CDC-Influenza Hospitalizations Admissions 3.07M 27.1% Weekly Hospitalized Case Counts
CinC 2012[77] originates from the PhysioNet 2012 Challenge and contains multivariate time series from ICU patients. Each record includes asynchronously sampled clinical variables such as blood pressure, heart rate, and oxygen saturation, with irregular measurement frequencies and observation patterns. The forecasting task focuses on predicting future physiological values and deterioration risk based on past observations.
CinC 2019[78] Derived from the PhysioNet 2019 Challenge, this dataset includes time series recorded during in-hospital stays, including variables like vital signs, lab test results, and intervention indicators. The data are highly sparse and irregular, reflecting realistic clinical workflows. We focus on continuous value prediction using sliding window evaluation, emphasizing temporal robustness.
CDC-IHA [^4] https://data.cdc.gov/Public-Health-Surveillance, published by the U.S. CDC, contains weekly aggregated hospital admission metrics related to respiratory diseases across U.S. jurisdictions. It includes counts for COVID-19, influenza, and RSV-related hospitalizations. The signals are inherently discrete and asynchronous across regions, with missing entries due to delayed or inconsistent reporting. The forecasting task involves predicting near-future hospitalization counts by location. We using below column in an weekly frequency: Total Patients Hospitalized with COVID-19, Total Patients Hospitalized with Influenza,Total Patients Hospitalized with RSV, Total ICU Patients Hospitalized with COVID-19,Total ICU Patients Hospitalized with Influenza,Total ICU Patients Hospitalized with RSV.
WHO FluNet [^5] https://www.who.int/tools/flunet These epidemiological datasets from the World Health Organization comprise country-level weekly flu case reports (FluNet).
MIT-BIH [79] This dataset provides annotated ECG waveform segments from patients with arrhythmias. The input is a regularly sampled 2-lead ECG sequence, and the forecasting task involves predicting future signal windows. We introduce synthetic missingness during evaluation to test robustness.
JHU COVID-19 Dataset [80] This dataset aggregates daily COVID-19 cases and deaths by region. While data are reported at regular intervals, inconsistencies and reporting delays motivate the use of masked evaluation. Forecasting focuses on short-term case count prediction. We using this dataset in a daily frequency
CDC-Illness [82] The CDC outpatient illness dataset tracks ILI (influenza-like illness) and other symptoms across U.S. reporting centers. Although data are uniformly weekly, we mask 30% of time points to simulate partial surveillance dropout.
Heart Rate This dataset, sourced from the Monash Time Series Extrinsic Regression Archive [81], contains photoplethysmography (PPG) sequences paired with continuous heart rate values as targets. Each instance represents a short PPG segment with the goal of predicting the corresponding heart rate, making it a benchmark for physiological signal regression tasks.
Appendix H Baselines
Time-MoE [28]: A sparsely activated decoder-only Transformer model that incorporates a Mixture-of-Experts (MoE) architecture with token-level routing. Trained on over 300 billion real-world time points spanning nine domains, Time-MoE demonstrates strong zero-shot forecasting performance, especially in long-range and multi-resolution prediction. It uses autoregressive decoding and sliding window inference with shared expert regularization.
Moirai [27]: A universal forecasting backbone that uses patch-wise tokenization and any-variate self-attention. It is trained on the LOTSA dataset comprising 27 billion time points and supports forecasting across arbitrary time steps, variable sets, and resolutions. Moirai also employs resolution-adaptive projection layers to accommodate different patch sizes during inference.
Moirai-MoE [30]: An extension of Moirai that integrates a token-level MoE module into the decoder block. This allows for frequency-specific specialization without relying on handcrafted signal partitions. Moirai-MoE achieves improved generalization across domains with limited added compute, and supports both token-aware routing and auxiliary balancing loss.
TimesFM [29]: A decoder-only Transformer developed by Google Research, pre-trained on 100 billion real-world time points from diverse sources including IoT, finance, and weather. It features autoregressive generation with fixed-length context windows and shows strong performance in forecasting across hundreds of benchmarks.
Chronos [32]: A family of probabilistic time series foundation models that transform numerical sequences into quantized token representations. Chronos leverages discrete latent spaces and causal language modeling for forecasting and sampling. It supports both point and probabilistic prediction.
Timer [85]: A generative time series language model trained on a diverse mix of real-world and synthetic data. It uses next-token prediction and masked time-step modeling for flexible downstream adaptation. While architecture and scale details are less publicly documented, it has demonstrated competitive zero-shot accuracy on common forecasting benchmarks.
ContiFormer [18]: A continuous-time Transformer model that combines Neural ODEs with attention. It uses a NeuralODE kernel for modeling value transitions between observations and a standard self-attention block for relational reasoning. It is trained end-to-end on irregularly sampled series.
T-PatchGNN [87]: A graph-based model that converts univariate time series into a graph of overlapping time patches, using GNNs to capture local and global dependencies. It is especially effective for sparse and low-signal series and supports continuous-time node embedding.
Neural-CDE [88]: A continuous-time model based on controlled differential equations. It encodes the trajectory of observed data using a Neural CDE solver and predicts future values through learned hidden dynamics. Particularly suited for datasets with asynchronous observations.
ODE-RNN [65]: Combines standard RNN encoders with ODE solvers to update latent states over irregular intervals. It supports interpolation between time steps and improves temporal continuity compared to discrete RNNs.
Sundial [31] introduces a family of scalable time series foundation models trained on the extensive TimeBench dataset, comprising one trillion time points. Utilizing the novel TimeFlow Loss, Sundial enables generative pretraining of Transformers without discrete tokenization, achieving state-of-the-art zero-shot performance across diverse forecasting benchmarks.
Moment [83] presents open-source Transformer-based models pre-trained on the Time-series Pile, a large and diverse collection of public time series data. Employing a masked time series prediction task, MOMENT demonstrates strong performance across forecasting, classification, and anomaly detection tasks, particularly in low-resource settings .
Lag-Llama [86] is a decoder-only Transformer model tailored for univariate probabilistic forecasting. By incorporating lagged inputs as covariates and pretraining on a diverse corpus of time series data, Lag-Llama exhibits robust zero-shot generalization and state-of-the-art performance upon fine-tuning on unseen datasets .
TimeGPT [84] is introduced as the first foundation model for time series analysis, capable of generating accurate predictions across diverse datasets without additional training. Leveraging advancements in deep learning, TimeGPT’s zero-shot inference outperforms traditional statistical, machine learning, and deep learning methods in both performance and efficiency .
Appendix I Evaluation Metrics
We evaluate forecasting performance using two widely adopted metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Both are computed over all predicted time steps and variables, averaged across evaluation windows.
Mean Absolute Error (MAE) quantifies the average absolute deviation between predictions $\hat{x}_{t}$ and ground truth $x_{t}$:
$\text{MAE}=\frac{1}{N}\sum_{t=1}^{N}|x_{t}-\hat{x}_{t}|,$ (23)
where $N$ denotes the number of valid (unmasked) prediction points. MAE is robust to outliers and provides a direct measure of average deviation.
Root Mean Squared Error (RMSE) emphasizes larger errors by squaring deviations before averaging:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{t=1}^{N}(x_{t}-\hat{x}_{t})^{2}}.$ (24)
It is more sensitive to large prediction errors and penalizes high-variance outputs.
Appendix J Detailed Performance Analysis of Out-of-Distribution Benchmarks
We proved detail performance of different baselines and our model.
Table 9: Zero-shot forecasting performance on out-of-distribution datasets. Reported values are predicted by zero-shot forecasting model pre-trained on general corpora.
Models Lag-llama TimeGPT Timer Momentsmall Momentbase Momentlarge Moirai-MoEsmall Moirai-MoEbase
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) 24 $1.643$ $1.385$ $1.668$ $1.463$ $1.899$ $1.769$ $3.231$ $2.035$ $3.218$ $2.025$ $3.204$ $2.019$ $1.443$ $1.137$ $1.636$ $1.321$
32 $1.702$ $1.433$ $2.090$ $1.755$ $1.898$ $1.721$ $2.985$ $1.849$ $2.978$ $1.857$ $2.949$ $1.479$ $1.777$ $1.432$ $1.942$ $1.595$
48 $1.813$ $1.528$ $2.592$ $2.198$ $1.898$ $1.667$ $2.787$ $1.738$ $2.764$ $1.726$ $2.734$ $1.711$ $2.374$ $1.927$ $2.523$ $2.095$
64 $1.895$ $1.605$ $2.683$ $2.245$ $1.907$ $1.657$ $2.861$ $1.784$ $2.796$ $1.732$ $2.781$ $1.731$ $2.347$ $1.906$ $2.476$ $1.595$
Avg. $1.764$ $1.488$ $2.258$ $1.915$ $1.901$ $1.704$ $2.966$ $1.852$ $2.939$ $1.835$ $2.917$ $1.735$ $1.982$ $1.600$ $2.144$ $1.652$
[…]
32 $0.185$ $0.148$ $0.175$ $0.141$ $0.219$ $0.184$ $0.407$ $0.216$ $0.405$ $0.214$ $0.402$ $0.211$ $0.185$ $0.152$ $0.185$ $0.151$
48 $0.253$ $0.192$ $0.291$ $0.227$ $0.330$ $0.278$ $0.407$ $0.224$ $0.406$ $0.223$ $0.402$ $0.218$ $0.231$ $0.183$ $0.231$ $0.183$
64 $0.268$ $0.203$ $0.321$ $0.256$ $0.292$ $0.230$ $0.401$ $0.216$ $0.402$ $0.217$ $0.396$ $0.210$ $0.252$ $0.198$ $0.253$ $0.198$
Avg. $0.217$ $0.169$ $0.231$ $0.185$ $0.255$ $0.213$ $0.417$ $0.229$ $0.416$ $0.228$ $0.413$ $0.224$ $0.208$ $0.167$ $0.208$ $0.167$
[…]
18 $5.193$ $3.968$ $5.208$ $3.989$ $4.908$ $3.727$ $12.884$ $4.435$ $12.941$ $4.404$ $12.894$ $4.416$ $5.751$ $4.600$ $5.868$ $4.731$
24 $7.643$ $5.528$ $7.886$ $5.535$ $7.670$ $5.601$ $22.026$ $5.921$ $21.651$ $5.934$ $21.735$ $5.887$ $7.852$ $5.715$ $8.167$ $6.077$
30 $6.935$ $5.162$ $7.206$ $5.245$ $6.881$ $5.104$ $20.436$ $5.593$ $20.008$ $5.524$ $20.363$ $5.511$ $7.852$ $5.715$ $7.906$ $6.001$
Avg. $6.531$ $4.846$ $6.654$ $4.860$ $6.424$ $4.857$ $17.803$ $5.307$ $17.631$ $5.288$ $17.689$ $5.260$ $7.099$ $5.405$ $7.639$ $5.862$
JH COVID-19 ($10^{2}$) 7 $2.115$ $0.825$ $0.160$ $0.156$ $3.141$ $3.141$ $2.667$ $0.404$ $2.662$ $0.403$ $2.672$ $0.404$ $0.467$ $0.356$ $0.487$ $0.357$
14 $3.052$ $1.214$ $0.990$ $0.817$ $2.268$ $2.135$ $3.795$ $0.642$ $3.826$ $0.644$ $3.811$ $0.641$ $0.547$ $0.396$ $0.522$ $0.389$
28 $3.874$ $1.588$ $1.279$ $0.892$ $2.415$ $2.118$ $4.096$ $0.786$ $4.143$ $0.795$ $4.048$ $0.779$ $0.924$ $0.661$ $0.916$ $0.674$
90 $5.342$ $2.101$ $5.087$ $4.459$ $2.765$ $1.917$ $1.749$ $0.378$ $1.756$ $0.377$ $1.724$ $0.373$ $3.871$ $1.486$ $75.64$ $11.34$
Avg. $3.596$ $1.432$ $1.879$ $1.580$ $2.647$ $2.328$ $3.077$ $0.553$ $3.097$ $0.554$ $3.064$ $0.549$ $1.452$ $0.725$ $19.391$ $3.190$
[…]
36 $1.702$ $1.321$ $2.729$ $1.582$ $2.316$ $1.751$ $1.467$ $0.988$ $1.466$ $0.986$ $1.464$ $0.983$ $1.995$ $1.667$ $2.033$ $1.701$
48 $1.914$ $1.493$ $2.337$ $1.227$ $2.051$ $1.549$ $1.669$ $1.134$ $1.663$ $1.128$ $1.662$ $1.128$ $2.158$ $1.801$ $2.119$ $1.765$
60 $2.118$ $1.659$ $2.075$ $0.991$ $1.914$ $1.493$ $1.594$ $1.111$ $1.590$ $1.110$ $1.589$ $1.106$ $2.218$ $1.859$ $2.195$ $1.842$
Avg. $1.780$ $1.366$ $2.011$ $1.077$ $1.882$ $1.492$ $1.570$ $1.056$ $1.566$ $1.054$ $1.566$ $1.052$ $2.001$ $1.678$ $1.983$ $1.664$
Table 10: Zero-shot forecasting performance on out-of-distribution datasets continued from last table. Reported values are predicted by zero-shot forecasting model pre-trained on general corpora.
Models Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge TimesFM
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) 24 $2.326$ $1.921$ $2.208$ $1.773$ $2.032$ $1.597$ $0.450$ $0.364$ $0.602$ $0.469$ $1.283$ $0.521$ $1.131$ $0.444$ $1.185$ $0.484$ $1.621$ $0.730$
32 $2.358$ $1.967$ $2.189$ $1.771$ $2.070$ $1.597$ $0.477$ $0.383$ $0.677$ $0.527$ $1.391$ $0.642$ $1.241$ $0.536$ $1.243$ $0.539$ $1.712$ $0.813$
48 $2.332$ $1.955$ $2.117$ $1.759$ $2.145$ $1.691$ $0.505$ $0.401$ $0.950$ $0.724$ $1.535$ $0.703$ $1.353$ $0.583$ $1.391$ $0.602$ $1.894$ $0.921$
64 $2.419$ $2.015$ $2.109$ $1.764$ $2.145$ $1.691$ $0.333$ $0.270$ $1.103$ $0.837$ $1.196$ $0.484$ $1.001$ $0.394$ $1.024$ $0.398$ $1.784$ $0.853$
Avg. $2.359$ $1.965$ $2.156$ $1.767$ $2.098$ $1.644$ $0.850$ $0.650$ $0.833$ $0.639$ $1.357$ $0.587$ $1.189$ $0.489$ $1.218$ $0.506$ $1.753$ $0.832$
[…]
32 $0.311$ $0.231$ $0.336$ $0.253$ $0.443$ $0.111$ $0.188$ $0.146$ $0.191$ $0.149$ $0.390$ $0.169$ $0.407$ $0.171$ $0.392$ $0.167$ $0.375$ $0.155$
48 $0.372$ $0.264$ $0.457$ $0.329$ $0.715$ $0.146$ $0.205$ $0.155$ $0.205$ $0.155$ $0.367$ $0.162$ $0.368$ $0.167$ $0.352$ $0.162$ $0.347$ $0.158$
64 $0.401$ $0.282$ $0.579$ $0.392$ $0.719$ $0.191$ $0.152$ $0.121$ $0.152$ $0.122$ $0.323$ $0.128$ $0.331$ $0.131$ $0.317$ $0.131$ $0.305$ $0.126$
Avg. $0.343$ $0.249$ $0.421$ $0.302$ $0.593$ $0.149$ $0.171$ $0.135$ $0.172$ $0.135$ $0.353$ $0.147$ $0.361$ $0.149$ $0.350$ $0.147$ $0.335$ $0.141$
[…]
18 $5.761$ $4.589$ $6.277$ $4.877$ $5.856$ $4.731$ $4.986$ $3.854$ $4.964$ $3.787$ $10.711$ $3.407$ $11.099$ $3.581$ $11.693$ $3.832$ $10.666$ $3.487$
24 $7.939$ $5.783$ $8.201$ $6.102$ $7.759$ $5.807$ $7.707$ $5.636$ $7.711$ $5.624$ $19.293$ $5.366$ $19.341$ $5.219$ $19.402$ $5.289$ $19.737$ $5.121$
30 $7.381$ $5.698$ $8.099$ $5.802$ $7.144$ $5.448$ $6.839$ $5.056$ $6.871$ $4.996$ $18.273$ $4.729$ $19.223$ $4.745$ $19.163$ $4.766$ $18.302$ $4.804$
Avg. $6.835$ $5.271$ $7.328$ $5.526$ $6.788$ $5.302$ $6.311$ $4.748$ $6.312$ $4.715$ $15.502$ $4.421$ $15.825$ $4.438$ $15.986$ $4.517$ $15.633$ $4.408$
JH COVID-19 ($10^{2}$) 7 $0.548$ $0.392$ $0.416$ $0.327$ $0.451$ $0.334$ $0.412$ $0.321$ $0.346$ $0.281$ $2.834$ $0.341$ $2.038$ $0.257$ $2.113$ $0.254$ $1.903$ $0.221$
14 $0.998$ $0.694$ $0.522$ $0.388$ $0.484$ $0.365$ $0.654$ $0.434$ $0.421$ $0.331$ $3.626$ $0.570$ $3.058$ $0.375$ $2.993$ $0.361$ $2.653$ $0.305$
28 $3.157$ $0.717$ $0.984$ $0.702$ $0.760$ $0.551$ $0.621$ $0.438$ $0.651$ $0.426$ $5.336$ $1.164$ $3.842$ $0.564$ $4.063$ $0.585$ $3.181$ $0.411$
90 $2.966$ $0.977$ $2.042$ $0.479$ $0.761$ $0.358$ $0.697$ $0.416$ $0.632$ $0.446$ $7.506$ $2.051$ $6.401$ $1.006$ $4.741$ $0.884$ $1.581$ $0.349$
Avg. $1.917$ $0.695$ $0.991$ $0.474$ $0.614$ $0.402$ $0.596$ $0.402$ $0.512$ $0.371$ $4.826$ $1.031$ $3.835$ $0.551$ $3.478$ $0.521$ $2.329$ $0.322$
[…]
36 $2.044$ $1.707$ $1.972$ $1.639$ $1.831$ $1.511$ $1.334$ $0.989$ $1.418$ $1.049$ $2.115$ $1.504$ $1.977$ $1.382$ $1.904$ $1.315$ $2.116$ $1.384$
48 $2.138$ $1.780$ $1.967$ $1.631$ $1.977$ $1.616$ $1.441$ $1.035$ $1.491$ $1.073$ $2.249$ $1.521$ $2.127$ $1.433$ $2.031$ $1.368$ $2.249$ $1.521$
60 $2.189$ $1.827$ $2.023$ $1.676$ $1.989$ $1.648$ $1.373$ $0.978$ $1.425$ $1.027$ $2.121$ $1.431$ $2.034$ $1.362$ $1.947$ $1.298$ $2.121$ $1.431$
Avg. $1.995$ $1.671$ $1.871$ $1.561$ $1.808$ $1.499$ $1.288$ $0.951$ $1.366$ $1.015$ $2.054$ $1.400$ $1.940$ $1.308$ $1.870$ $1.252$ $2.034$ $1.333$
Table 11: Zero-shot forecasting performance on out-of-distribution datasets. Reported values are predicted by zero-shot forecasting model pre-trained on medical corpora.
Models Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Heart Rate ($10^{-1}$) 24 $1.555$ $1.254$ $1.385$ $1.082$ $1.203$ $0.956$ $0.753$ $0.616$ $0.682$ $0.562$ $0.933$ $0.451$ $0.872$ $0.402$ $0.815$ $0.381$
32 $2.418$ $2.002$ $2.283$ $1.850$ $1.849$ $1.464$ $0.647$ $0.523$ $0.614$ $0.495$ $1.014$ $0.582$ $0.935$ $0.521$ $0.879$ $0.490$
48 $2.375$ $1.967$ $2.269$ $1.842$ $1.960$ $1.524$ $0.714$ $0.567$ $0.656$ $0.519$ $1.255$ $0.703$ $1.153$ $0.633$ $1.091$ $0.602$
64 $1.843$ $1.515$ $1.692$ $1.370$ $1.387$ $1.108$ $0.477$ $0.390$ $0.461$ $0.375$ $0.996$ $0.484$ $0.901$ $0.394$ $0.824$ $0.358$
Avg. $2.047$ $1.685$ $1.907$ $1.536$ $1.601$ $1.263$ $0.648$ $0.524$ $0.603$ $0.488$ $1.049$ $0.555$ $0.965$ $0.488$ $0.902$ $0.458$
[…]
32 $0.251$ $0.209$ $0.261$ $0.203$ $0.247$ $0.196$ $0.173$ $0.133$ $0.166$ $0.128$ $0.350$ $0.159$ $0.367$ $0.161$ $0.352$ $0.147$
48 $0.257$ $0.210$ $0.456$ $0.253$ $0.272$ $0.214$ $0.136$ $0.101$ $0.139$ $0.104$ $0.327$ $0.142$ $0.328$ $0.147$ $0.312$ $0.132$
64 $0.232$ $0.203$ $0.199$ $0.158$ $0.191$ $0.155$ $0.083$ $0.066$ $0.087$ $0.070$ $0.283$ $0.118$ $0.291$ $0.121$ $0.277$ $0.111$
Avg. $0.239$ $0.204$ $0.274$ $0.190$ $0.219$ $0.176$ $0.132$ $0.103$ $0.131$ $0.101$ $0.311$ $0.137$ $0.320$ $0.139$ $0.306$ $0.127$
[…]
18 $5.656$ $4.639$ $5.754$ $4.553$ $5.426$ $4.262$ $5.054$ $3.864$ $4.987$ $3.749$ $9.711$ $3.307$ $10.099$ $3.481$ $10.693$ $3.732$
24 $7.775$ $5.948$ $7.958$ $5.982$ $7.708$ $5.669$ $7.668$ $5.602$ $7.655$ $5.571$ $18.293$ $5.266$ $18.341$ $5.119$ $18.402$ $5.189$
30 $7.086$ $5.500$ $7.420$ $5.713$ $7.043$ $5.310$ $6.924$ $5.030$ $6.851$ $5.000$ $17.273$ $4.629$ $18.223$ $4.645$ $18.163$ $4.666$
Avg. $6.690$ $5.310$ $6.934$ $5.376$ $6.696$ $5.114$ $6.327$ $4.698$ $6.299$ $4.666$ $14.502$ $4.321$ $14.825$ $4.338$ $14.986$ $4.417$
JH COVID-19 ($10^{2}$) 7 $0.373$ $0.311$ $0.599$ $0.442$ $0.385$ $0.201$ $0.340$ $0.272$ $0.329$ $0.258$ $2.634$ $0.301$ $2.038$ $0.257$ $1.913$ $0.204$
14 $0.531$ $0.415$ $0.537$ $0.437$ $0.719$ $0.277$ $0.477$ $0.354$ $0.425$ $0.334$ $3.426$ $0.470$ $2.958$ $0.315$ $2.893$ $0.301$
28 $0.707$ $0.515$ $0.586$ $0.497$ $0.200$ $0.527$ $0.689$ $0.473$ $0.712$ $0.484$ $4.336$ $1.064$ $3.742$ $0.514$ $3.863$ $0.535$
90 $0.705$ $0.549$ $0.753$ $0.535$ $1.942$ $0.343$ $0.531$ $0.315$ $0.602$ $0.372$ $6.506$ $1.951$ $5.401$ $0.956$ $4.641$ $0.834$
Avg. $0.579$ $0.448$ $0.619$ $0.478$ $0.812$ $0.337$ $0.509$ $0.353$ $0.517$ $0.362$ $4.225$ $0.947$ $3.535$ $0.510$ $3.328$ $0.469$
[…]
36 $1.581$ $1.327$ $1.457$ $1.204$ $1.546$ $1.465$ $1.221$ $0.939$ $1.245$ $0.960$ $1.911$ $1.188$ $1.822$ $1.165$ $1.704$ $1.115$
48 $1.556$ $1.279$ $1.474$ $1.186$ $1.049$ $1.345$ $1.318$ $0.993$ $1.333$ $1.006$ $1.692$ $1.200$ $1.654$ $1.166$ $1.531$ $1.068$
60 $1.499$ $1.232$ $1.454$ $1.174$ $0.843$ $1.208$ $1.255$ $0.943$ $1.265$ $0.948$ $1.655$ $1.123$ $1.571$ $1.062$ $1.447$ $0.998$
Avg. $1.528$ $1.289$ $1.435$ $1.191$ $1.501$ $1.229$ $1.188$ $0.915$ $1.201$ $0.927$ $1.705$ $1.127$ $1.639$ $1.081$ $1.547$ $1.028$
Table 12: Zero-shot Forecasting on out-of-distribution datasets continued from last table.
Models MIRAsmall MIRAbase MIRAlarge Contiformer T-PatchGNN ODE-RNN Neural-CDE
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
[…]
32 $1.144$ $0.964$ $1.089$ $0.965$ $0.941$ $0.815$ $0.784$ $0.637$ $0.613$ $0.492$ $0.943$ $0.677$ $0.651$ $0.584$
48 $1.208$ $1.012$ $1.116$ $1.048$ $0.984$ $0.847$ $0.775$ $0.623$ $0.626$ $0.494$ $0.967$ $0.691$ $0.692$ $0.594$
64 $1.315$ $1.114$ $1.229$ $0.914$ $1.086$ $0.941$ $0.799$ $0.635$ $0.616$ $0.501$ $0.989$ $0.719$ $0.723$ $0.609$
Avg. $1.195$ $1.001$ $1.123$ $0.950$ $0.986$ $0.845$ $0.774$ $0.633$ $0.627$ $0.497$ $0.945$ $0.683$ $0.671$ $0.587$
[…]
32 $0.261$ $0.181$ $0.237$ $0.163$ $0.128$ $0.103$ $0.453$ $0.354$ $0.675$ $0.606$ $0.882$ $0.623$ $0.242$ $0.196$
48 $0.322$ $0.214$ $0.290$ $0.193$ $0.104$ $0.087$ $0.454$ $0.355$ $0.717$ $0.649$ $0.884$ $0.625$ $0.243$ $0.197$
64 $0.351$ $0.232$ $0.319$ $0.205$ $0.070$ $0.058$ $0.452$ $0.353$ $0.691$ $0.595$ $0.881$ $0.622$ $0.242$ $0.195$
Avg. $0.293$ $0.198$ $0.263$ $0.176$ $0.107$ $0.088$ $0.453$ $0.354$ $0.705$ $0.627$ $0.882$ $0.623$ $0.242$ $0.196$
[…]
18 $5.943$ $4.753$ $5.615$ $4.427$ $5.296$ $4.215$ $5.501$ $4.123$ $8.824$ $7.383$ $10.073$ $9.036$ $7.823$ $6.717$
24 $6.143$ $4.710$ $5.931$ $4.459$ $5.630$ $4.315$ $5.535$ $4.137$ $8.388$ $7.005$ $10.116$ $9.089$ $7.943$ $6.774$
30 $6.475$ $5.048$ $6.258$ $4.912$ $5.924$ $4.571$ $5.490$ $4.159$ $8.778$ $7.445$ $10.214$ $9.131$ $8.022$ $6.896$
Avg. $5.976$ $4.684$ $5.729$ $4.502$ $5.426$ $4.212$ $5.51$ $4.133$ $9.522$ $7.974$ $10.068$ $9.052$ $7.892$ $6.766$
[…]
14 $0.395$ $0.342$ $0.382$ $0.330$ $0.354$ $0.302$ $0.313$ $0.287$ $0.213$ $0.174$ $0.416$ $0.321$ $0.511$ $0.486$
28 $0.432$ $0.381$ $0.418$ $0.365$ $0.388$ $0.337$ $0.341$ $0.304$ $0.585$ $0.443$ $0.432$ $0.343$ $0.577$ $0.529$
90 $0.461$ $0.408$ $0.451$ $0.397$ $0.423$ $0.372$ $0.374$ $0.351$ $0.551$ $0.497$ $0.447$ $0.357$ $0.604$ $0.535$
Avg. $0.407$ $0.355$ $0.395$ $0.342$ $0.366$ $0.315$ $0.323$ $0.297$ $0.350$ $0.291$ $0.424$ $0.331$ $0.545$ $0.503$
[…]
36 $1.303$ $1.094$ $1.225$ $1.043$ $1.135$ $0.962$ $0.388$ $0.222$ $0.195$ $0.144$ $0.406$ $0.262$ $0.423$ $0.314$
48 $1.346$ $1.101$ $1.278$ $1.054$ $1.155$ $0.973$ $0.393$ $0.225$ $0.195$ $0.143$ $0.411$ $0.265$ $0.422$ $0.314$
60 $1.415$ $1.182$ $1.333$ $1.117$ $1.209$ $1.012$ $0.407$ $0.231$ $0.195$ $0.143$ $0.418$ $0.268$ $0.426$ $0.317$
Avg. $1.294$ $1.077$ $1.218$ $1.024$ $1.113$ $0.943$ $0.391$ $0.224$ $0.195$ $0.143$ $0.410$ $0.264$ $0.423$ $0.314$
Appendix K Detailed Performance Analysis of In-Distribution Benchmarks
We proved detail performance of all the model pre-training on the medical corpus.
Table 13: Detail Performance on zero-shot forecasting performance at in-distribution setting.
Models MIRAsmall MIRAbase MIRAlarge Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
SleepEDF ($10^{2}$) 24 $0.210$ $0.177$ $0.191$ $0.160$ $0.185$ $0.154$ $0.305$ $0.277$ $0.394$ $0.164$ $0.218$ $0.241$ $0.226$ $0.189$ $0.243$ $0.203$ $0.359$ $0.167$ $0.357$ $0.165$ $0.357$ $0.165$
36 $0.219$ $0.184$ $0.199$ $0.165$ $0.193$ $0.158$ $0.308$ $0.269$ $0.554$ $0.181$ $0.271$ $0.312$ $0.237$ $0.196$ $0.261$ $0.215$ $0.445$ $0.207$ $0.450$ $0.210$ $0.448$ $0.209$
48 $0.232$ $0.191$ $0.211$ $0.172$ $0.204$ $0.166$ $0.302$ $0.259$ $0.509$ $0.169$ $0.348$ $0.393$ $0.245$ $0.202$ $0.257$ $0.212$ $0.456$ $0.214$ $0.464$ $0.217$ $0.464$ $0.218$
60 $0.197$ $0.168$ $0.178$ $0.151$ $0.172$ $0.145$ $0.284$ $0.251$ $1.214$ $0.393$ $0.378$ $0.346$ $0.205$ $0.172$ $0.216$ $0.181$ $0.383$ $0.178$ $0.384$ $0.179$ $0.381$ $0.178$
Avg. $0.215$ $0.180$ $0.195$ $0.162$ $0.189$ $0.156$ $0.301$ $0.264$ $0.668$ $0.227$ $0.304$ $0.323$ $0.228$ $0.191$ $0.244$ $0.203$ $0.411$ $0.192$ $0.414$ $0.193$ $0.413$ $0.193$
[…]
36 $0.142$ $0.106$ $0.122$ $0.091$ $0.116$ $0.087$ $0.181$ $0.123$ $0.209$ $0.082$ $0.303$ $0.075$ $0.114$ $0.065$ $0.112$ $0.063$ $0.224$ $0.100$ $0.225$ $0.104$ $0.224$ $0.104$
48 $0.153$ $0.114$ $0.132$ $0.099$ $0.126$ $0.095$ $0.192$ $0.131$ $0.361$ $0.104$ $0.70$ $0.131$ $0.121$ $0.064$ $0.121$ $0.063$ $0.226$ $0.102$ $0.226$ $0.104$ $0.223$ $0.104$
60 $0.164$ $0.123$ $0.143$ $0.107$ $0.136$ $0.102$ $0.176$ $0.130$ $0.280$ $0.113$ $0.359$ $0.101$ $0.091$ $0.055$ $0.091$ $0.068$ $0.218$ $0.092$ $0.238$ $0.099$ $0.226$ $0.096$
Avg. $0.147$ $0.110$ $0.127$ $0.095$ $0.121$ $0.091$ $0.177$ $0.125$ $0.270$ $0.098$ $0.416$ $0.099$ $0.110$ $0.063$ $0.109$ $0.066$ $0.228$ $0.100$ $0.234$ $0.104$ $0.229$ $0.103$
[…]
36 $0.120$ $0.101$ $0.101$ $0.084$ $0.096$ $0.080$ $0.167$ $0.144$ $0.310$ $0.109$ $0.148$ $0.142$ $0.108$ $0.081$ $0.107$ $0.079$ $0.171$ $0.087$ $0.168$ $0.088$ $0.166$ $0.088$
48 $0.132$ $0.110$ $0.112$ $0.093$ $0.107$ $0.088$ $0.179$ $0.152$ $0.187$ $0.142$ $0.232$ $0.156$ $0.119$ $0.104$ $0.116$ $0.096$ $0.193$ $0.098$ $0.195$ $0.100$ $0.193$ $0.102$
60 $0.143$ $0.119$ $0.123$ $0.102$ $0.117$ $0.097$ $0.162$ $0.143$ $0.278$ $0.307$ $0.187$ $0.167$ $0.093$ $0.061$ $0.091$ $0.059$ $0.127$ $0.066$ $0.131$ $0.069$ $0.123$ $0.068$
Avg. $0.126$ $0.106$ $0.107$ $0.089$ $0.102$ $0.084$ $0.163$ $0.141$ $0.256$ $0.164$ $0.172$ $0.138$ $0.105$ $0.081$ $0.103$ $0.078$ $0.153$ $0.079$ $0.154$ $0.080$ $0.151$ $0.080$
[…]
36 $0.105$ $0.089$ $0.090$ $0.065$ $0.082$ $0.056$ $0.272$ $0.229$ $0.283$ $0.275$ $0.353$ $0.165$ $0.104$ $0.082$ $0.103$ $0.079$ $0.366$ $0.171$ $0.355$ $0.165$ $0.353$ $0.164$
48 $0.117$ $0.099$ $0.096$ $0.074$ $0.093$ $0.066$ $0.269$ $0.216$ $0.304$ $0.318$ $0.374$ $0.174$ $0.091$ $0.064$ $0.088$ $0.062$ $0.356$ $0.155$ $0.372$ $0.173$ $0.374$ $0.174$
60 $0.129$ $0.108$ $0.108$ $0.081$ $0.104$ $0.073$ $0.236$ $0.210$ $0.367$ $0.375$ $0.264$ $0.117$ $0.052$ $0.040$ $0.047$ $0.038$ $0.229$ $0.096$ $0.257$ $0.114$ $0.264$ $0.117$
Avg. $0.111$ $0.094$ $0.091$ $0.069$ $0.081$ $0.061$ $0.259$ $0.223$ $0.300$ $0.291$ $0.319$ $0.143$ $0.084$ $0.064$ $0.082$ $0.062$ $0.309$ $0.134$ $0.317$ $0.142$ $0.319$ $0.143$
[…]
36 $0.155$ $0.129$ $0.137$ $0.112$ $0.130$ $0.107$ $0.184$ $0.161$ $0.192$ $0.174$ $0.169$ $0.149$ $0.156$ $0.129$ $0.143$ $0.118$ $0.191$ $0.125$ $0.193$ $0.126$ $0.192$ $0.125$
48 $0.167$ $0.139$ $0.148$ $0.122$ $0.141$ $0.116$ $0.185$ $0.161$ $0.222$ $0.177$ $0.208$ $0.153$ $0.157$ $0.131$ $0.150$ $0.124$ $0.219$ $0.137$ $0.217$ $0.136$ $0.217$ $0.135$
60 $0.151$ $0.126$ $0.133$ $0.110$ $0.126$ $0.103$ $0.169$ $0.151$ $0.160$ $0.194$ $0.152$ $0.136$ $0.129$ $0.108$ $0.124$ $0.104$ $0.171$ $0.110$ $0.169$ $0.109$ $0.167$ $0.108$
Avg. $0.154$ $0.129$ $0.136$ $0.112$ $0.129$ $0.106$ $0.177$ $0.157$ $0.190$ $0.181$ $0.169$ $0.155$ $0.148$ $0.124$ $0.141$ $0.116$ $0.184$ $0.119$ $0.183$ $0.118$ $0.182$ $0.117$